OpenStudy (anonymous):

3=2+2/z+2 i know the answer is 0 but i need to know how in steps

OpenStudy (saifoo.khan):

The answer isnt 0/

OpenStudy (mertsj):

Cross multiply and get 3z + 6 = 4 3z = -2 z = -2/3

OpenStudy (saifoo.khan):

wait.

OpenStudy (anonymous):

2 over z+2 i couldnt do it like that , that is why it look like that

OpenStudy (saifoo.khan):

\[3=2+\frac2z+2\]\[3 = \frac 2x + 4\]\[-1=\frac2x\]\[x=-\frac21\]\[x=-2\]

OpenStudy (mertsj):

Then multiply each term by z+2 and get: 3z + 6 =2z+4 +2 3z = 6 = 2z +6 Subtract 2z from both sides and get : z + 6 = 6 Subtract 6 from both sides and get z= 0

OpenStudy (mertsj):

She means 3 = 2 + 2/(z+2)

OpenStudy (anonymous):

yes thanks

OpenStudy (saifoo.khan):

justdat, can u use the draw tool to make it?

OpenStudy (anonymous):

lol i am need to this didnt know

OpenStudy (mertsj):